OpenStudy (anonymous):

ill give a medal . And its multiple choice. Annika mixed red and blue color candies. Annika bought 8 pounds of red color candies for $22. The cost per pound for blue color candies is 48% more than the cost per pound red color candies. Annika bought enough blue color candies that, when she mixed them with the red color candies, the mixture had a value of $3.21 per pound. Approximately what percent of the mixture, by weight, was blue color candies? 47% 65% 28% 35%

OpenStudy (anonymous):

what a drag of a word problem i am sure we can do it though

OpenStudy (anonymous):

8 pound for $22 what is the price per pound?

OpenStudy (anonymous):

2.75

OpenStudy (anonymous):

ok good now lets get the price of the blue candies. what is 48% more than \(2.75\) ?

OpenStudy (anonymous):

i dont know ??

OpenStudy (anonymous):

oh ok lets find that there are two ways to do it one is to compute 48% of \(2.75\) via \(27.5\times .4\) and add it to \(2.75\) or an easier method with a calculator is to multiply \[2.75\times 1.48\]

OpenStudy (anonymous):

lets multiply http://www.wolframalpha.com/input/?i=2.75*1.48

OpenStudy (anonymous):

thats 4.07

OpenStudy (anonymous):

ok now we are in the home stretch

OpenStudy (anonymous):

right..

OpenStudy (anonymous):

Whats a home stretch ?

OpenStudy (anonymous):

i guess it is a horse racing analogy coming down to the finish line, aka almost done

OpenStudy (anonymous):

Oh haha okay .

OpenStudy (anonymous):

we got 8 pounds in red, lets say \(x\) in blue, for a total of \[8+x\] and a value \(3.21(8+x)\) which we also know is \[22+4.07x\]

OpenStudy (anonymous):

set \[3.21(8+x)=22+4.07x\] and solve for \(x\)

OpenStudy (anonymous):

or clear the decimal and solve \[321(8+x)=2200+407x\]

OpenStudy (anonymous):

it says "approximate" so lets call it \(4.3\) pounds http://www.wolframalpha.com/input/?i=321%288%2Bx%29%3D2200%2B407x

OpenStudy (anonymous):

and finally to answer as a percent compute \[\frac{4.3}{8+4.3}\]

OpenStudy (anonymous):

i hope one answer is \(35\%\)

OpenStudy (anonymous):

it is ! omg . thank you but where did 35 come from i got like 427. something percent!!

OpenStudy (anonymous):

did you compute \[4.3\div 12.3\]?

OpenStudy (anonymous):

oh no . My bad

OpenStudy (anonymous):

the total number of pound approx is \(8+4.3=12.3\) and the total blue was \(4.3\) the the percent is \(4.3\div 12.3\) after you make that decimal in to a percent by moving the decimal place over i just did this http://www.wolframalpha.com/input/?i=4.3%2F%288%2B4.3%29

OpenStudy (anonymous):

clear now though right?

OpenStudy (anonymous):

Yes . thank you so much .

OpenStudy (anonymous):

yw, my pleasure